OpenStudy (anonymous):

Find (f^-1)'(a) for the function f, given a=-11 f(x)=1/27(x^5-2x^3) Please help! I'm not sure how to figure this out at all.

myininaya (myininaya):

\[(f^{-1})'(a)=\frac{1}{f'(f^{-1}(a))}\]

myininaya (myininaya):

so if a=-11 then \[f^{-1}(-11)=? => f(?)=11\]

myininaya (myininaya):

what can we plug into the function that will give us 11 \[\frac{1}{27}(x^5-2x^3)=-11\]

myininaya (myininaya):

by they way that is -11 (not 11---this is a type-0 above)

OpenStudy (anonymous):

Thank you! this really helped... i think ill be able to solve it on my own